OpenStudy (anonymous):

XYZ is a triangle with side lengths in the ratio 2 : 3 : 4. The perimeter of XYZ is 27 yards. Find the length of the longest side. A. 12 yards B. 3 yards C. 6 yards D. 9 yards

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

2x+3x+4x=27 9x=27 x=27/9 x=3 the longest side is 3*2=6 the answer is C Click good answer .if I deserve a medal,TY

OpenStudy (anonymous):

So when you get 3 , why do you multiply it by 2 ?

OpenStudy (anonymous):

in the ratio 2 : 3 : 4. 2 is longest side ,than multiply it by 2

OpenStudy (anonymous):

Ohh ok thanks ! : )